• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Lindgren

    Log entry time 16:28:46 on July 3,2010

    Entry number 329682

    keyword=Shift Summary

    8:00 Day shift beginning
    R. Lindgren SL
    Victoria Owens TO
    Jason Bono 3rd
    Jim Boyce 4th

    Hall in Control access
    Bogdan, Albert, Sergey, amd Samvel in the Hall troublshooting
    8:40 Bob Michaels entered the Hall

    9:50 Hall A in beam permit
    Waiting for 5 ua beam to start beam position test
    10:33 Beam in Hall 5 ua checking Raster
    10:40 Harp scan by MCC shows beam spot 250 x 250
    10:50 Run spot++ to determine position of beam
    12:30 Corrector magnet keeps tripping
    MCC trying to repair it.
    13:10 Still waiting for MCC
    13:15 Beam back from magnet power supply thermal repair bypass
    13:30 Changing target to lead diamond number 2
    14:10 Hall A Escorted Access to investigate the "flashing light", when the septum is on. When the septum is off, the light is on continously. Somebody probably left a flouresecent light there.
    We will also turn put the scattering chamber light on remote control
    so we can turn the light down to see the beam on the BeO target.


    14:40 Radcon says to leave the portable light in the septum. It has been there for weeks. Ed Folts was also called but has not returned our message yet. We are now back in beam permit and septum current is back up. MCC is continuing set up for max allowed current using ion gauge as vacuum limit. Waiting for beam.

    15:55 Beam back 15 ua CW target lead number 2

    Run Number Target Raster Position(y=0) Spot++ T1 Rate
    1525 Empty 4x4
    1526 Empty 12x8/6x4
    1527 Empty Same
    1528 Empty Same 2 KH
    1531 C hole Same BPM x = -1 Hitting frame 5 KH
    1532 Taking cosmics
    1533 C hole Same BPM x = +1 Missing frame
    1534 C hole Same BPM x = +2 No change
    1535 C hole Same BPM x = +3 Not much change 10 KH
    1536 C hole Same BPM x = +3.7 vac gauge tripping
    1537 C hole Same BPM x= -2 No hole yet 4.3 KH
    1538 C hole Same BPM x= -3 No hole yet 3.6 KH
    1539 C hole Same BPM x= -4 No hole yet (low?)1.7 KH
    1540 C hole Same BPM x= -5 No hole yet 3.1 KH
    1541 C hole Same BPM x= -6 No hole yet 3.0 KH
    No Run C hole Same BPM x= -4 No Hole 3.4 KH 1542 C hole Same BPM x= -1 No hole yet 4.4 KH
    1546 C hole Same BPM x= -1,y=+1 No hole yet 4.9 KH 1547 C hole Same BPM x= -1,y=-2 No hole yet 4.0 KH

    1554 Lead hole 2 Same BPM x= 0,y= 0 see hole 2.2 MH
    1555 Lead hole 2 Same BPM x= 0,y= +1 see hole move 2.2 MH
    1556 Lead Hole 2 Same BPM x= 0,y= -1 see hole move 2.5 MH
    1557 Lead Hole 2 Same BPM x= -2,y= 0 see hole move 2.5 MH
    1558 Lead Hole 2 Same BPM x= +2,y= 0 see hole move 2.5 MH